# Beamforming and Resource Allocation for Delay Optimization in RIS-Assisted OFDM Systems

### v3

@@ III-F Computational Complexity
The computational complexity of PPO-$\Theta$ is $\mathcal{O}(KNN_{t}(2M+1))$, while that of PPO-$N$ is $\mathcal{O}(KNN_{t})$. On the other hand, the active beamforming operation incurs a complexity of $\mathcal{O}(NMN_{t}+N)$. As a result, the overall computational complexity of the proposed algorithm is $\mathcal{O}(KNN_{t}(2M+1)+KNN_{t}+NMN_{t}+N)$.
TABLE I: Hyper-Parameters
@@
PPO-$\Theta$ replay buffer size 1000
PPO-N replay buffer size 1000
@@ IV Numerical Results
In this section, a comprehensive evaluation of the performance of the proposed DRL-based algorithm is conducted. Various simulation scenarios and baseline comparisons are considered to demonstrate its effectiveness in optimizing system performance.
@@ IV-A Simulation Settings
The packet capacity $L$ is set to 512 bits. In each episode, the communication duration is set to $T_{\text{sum}}=$ 1000 time slots, with each time slot lasting 1 ms. Since the position of user remains nearly fixed within 1 second, we assume that the user positions are fixed throughout each episode [^1] Note that, to accelerate the convergence of the algorithm, the user positions are randomly generated at each step within an episode during the training process.. The packet arrival rate for each user is set to $\lambda=9.5$. Throughout this work, unless stated otherwise, identical arrival rates are assumed for all users, i.e., $\lambda=\lambda_{1}=\lambda_{2}=\cdots=\lambda_{K}$. The detailed hyperparameters are presented in Table I.
Figure 5: Return for each training episode with increasing iterations.
@@ IV-B Baseline Methods
To reflect the performance of our proposed method, the following baseline methods are considered as follow.
  • Max sum rate: The algorithm proposed in [34] and [35] are respectively employed for subcarrier allocation and RIS reflection shift optimization, while MRT is used for beamforming at the BS.
  • Max minimum rate: This algorithm is the hybrid DRL algorithm with the reward to be $\min R_{k}[t]$ reward, as described in Section III-E, and it has not undergone transfer learning.
  • Random: The RIS reflection shift and subcarrier allocation are set randomly. The active beamforming at BS is optimized using the MRT.
@@
Figure 7: Average delay performance as packet arrival rate $\lambda$ varies.
@@ IV-C Convergence and Validity Analysis
To validate the effectiveness and convergence of the proposed method, experiments are conducted. Fig. 5 shows the cumulative reward for each training episode. The figure demonstrates a progressive increase in the cumulative reward throughout the training process, indicating the effectiveness of our proposed method. Besides, the performance gradually stabilizes in around 150 training episodes, despite some fluctuations due to channel variations and changes of packets in buffers.
@@ IV-D Latency Performance Analysis
Fig. 6 illustrates the average delay performance under varying packet arrival rates $\lambda$. The proposed algorithm consistently achieves the lowest average delay, demonstrating its superior capability in minimizing latency. In contrast, the “without RIS” scheme results in the highest delay, emphasizing the critical role of RIS in improving delay performance. Although the “max sum rate” method performs better than baseline schemes such as “random” and “without RIS”, its overall delay performance is clearly inferior to that of the proposed method. This is because it prioritizes the maximization of system throughput without considering user fairness or queue stability, which leads to higher delays for certain users. The “random” method yields persistently high delays, further underlining the importance of intelligent and delay-aware resource allocation. Moreover, as the $\lambda$ increases, the number of packets arriving at the buffer within each time slot grows accordingly, leading to aggravated buffer backlog and consequently resulting in an incremental increase in the average delay.
(a) The backlogged packets in each buffer of the proposed method.
(b) The backlogged packets in each buffer of the max minimum rate method.
(c) Transmission rates of the proposed method.
(d) Transmission rates of the max minimum rate method.
Figure 8: The change of the backlogged packets and transmission rates in each buffer over time step of the proposed method and max minimum rate method in brustiness scenarios.
(a) The backlogged packets in each buffer of the proposed method.
(b) The backlogged packets in each buffer of the max minimum rate method.
(c) Transmission rates of the proposed method.
(d) Transmission rates of the max minimum rate method.
Figure 9: The change of the backlogged packets and transmission rates in each buffer over time step of the proposed method and max minimum rate method in different traffic flows scenarios.
Figure 10: Impact of reflection elements $M$ on delay performance.
@@ IV-E Latency Analysis in Brustiness Scenarios
To evaluate the proposed algorithm under bursty traffic, we simulate sudden data arrivals by injecting 300 packets into buffers 1, 2, and 3 at time slots 100, 400, and 700, respectively. Fig. 7 presents a comparison of the delay performance between the proposed algorithm and the “max minimum rate” algorithm under different scenarios. It can be observed that, regardless of whether burstiness is present, the proposed algorithm consistently demonstrates superior performance, effectively validating its effectiveness, in handling burstiness.
To further explain this advantage, we analyze a specific case by examining buffer backlogs and transmission rates during the first 300 time slots. As shown in Fig. 8LABEL:sub@Fig:packagePropsed, the proposed algorithm effectively reduces backlog by dynamically adjusting resource allocation in response to bursty arrivals. In contrast, Fig. 8LABEL:sub@Fig:packageMaxMin shows persistent congestion in buffer 3 under the “max minimum rate” approach, indicating inefficient allocation. Fig. 8LABEL:sub@rateProposed further shows that the proposed method prioritizes users with heavier backlogs, rapidly increasing their transmission rates, such as user 3, leading to quick backlog clearance. Conversely, Fig. 8LABEL:sub@rateMaxMin reveals that the baseline allocates resources more rigidly, ignoring buffer states and failing to relieve congestion effectively, which results in higher delay.
@@ IV-F Impact of Reflection Elements $M$
Fig. 10 illustrates the impact of the number of RIS reflection elements on the average delay performance under varying arrival rates $\lambda$. As observed, when $M=36$, the average delay increases sharply once $\lambda$ exceeds 9.0, suggesting that the system struggles to accommodate high traffic loads. With $M=64$, the delay grows more moderately but still experiences a considerable rise at elevated arrival rates. In contrast, when $M=100$, the delay remains consistently low across all values of $\lambda$, demonstrating $M=100$ is sufficient enough to support the considered service load with $\lambda$ less than 10. This improvement is primarily attributed to the enhanced passive beamforming gain resulting from the increased number of RIS reflection elements, thereby improving the effective data rate. As a result, the system is able to accommodate higher traffic loads while maintaining low-latency communication performance.
@@ IV-G Latency Analysis under Different Traffic Flows
Figure 11: Average delay performance as distinct packet arrival rate $\lambda$ varies.
@@
To further explain this advantage, we analyze a specific case by examining buffer backlogs and transmission rates during the first 300 time slots. In Fig. 9LABEL:sub@Fig:packagePropsed, the proposed algorithm keeps buffer backlogs balanced despite differing arrival rates, showing its fairness. Upon the burst at time slot 100, it quickly reallocates resources to clear buffer 3. In contrast, Fig. 9LABEL:sub@Fig:packageMaxMin shows the “max minimum rate” approach leads to severe congestion in buffer 1 due to static rate-based allocation. Further, Fig. 9LABEL:sub@rateProposed reveals the proposed method allocates more resources to high arrival rate users and dynamically adapts to queue states. Before time slot 100, user 3 has a lower rate due to low arrival, but receives rapid support when the burst occurs. Conversely, Fig. 9LABEL:sub@rateMaxMin shows that the “max minimum rate” method rigidly enforces rate fairness, ignoring buffer states, resulting in inefficient resource use and longer delays.
TABLE II: The robustness of proposed approach on average delay.
@@
colspec = —X[c]—X[c]—X[c]—X[c]—X[c]—, hlines, vlines Algorithm & Delay taps ($L_{0}=6$, $L_{1}=3$, $L_{3}=4$) Distribution ($R=15$ m, $D_{3}=3$ m) Rician factor ($k_{\mathrm{BR}}=3$ dB, $k_{\mathrm{RU}}=4$ dB)
@@ IV-H Jitter Analysis
$\frac{1}{K}\sum_{k=1}^{K}\sqrt{\frac{1}{m(k)}\sum_{g=1}^{m(k)}\bigg{(}T_{D}^{(k)}(g)-\frac{1}{m(k)}\sum_{g=1}^{m(k)}T_{D}^{(k)}(g)\bigg{)}^{2}},$ (27)
As illustrated in Fig. 12, the proposed algorithm consistently achieves the lowest jitter across all values of $\lambda$, demonstrating its superior capability in balancing resource allocation and maintaining latency stability. In contrast, the “max-min rate” scheme yields slightly higher jitter, while the “max sum rate” approach, although effective in optimizing overall throughput, results in significantly increased jitter. The “random” allocation and “without RIS” schemes exhibit the worst jitter performance, underscoring the effectiveness of both learning-based optimization and RIS assistance in enhancing system delay stability. Moreover, as the arrival rate $\lambda$ increases, jitter also exhibits a rising trend, indicating that the system becomes increasingly unstable under higher traffic loads and struggles to maintain consistent delay performance.
Figure 12: Jitter performance as arrival rate $\lambda$ varies.
@@ IV-I Robustness Validation
Table II presents a robustness evaluation of the proposed algorithm in terms of average delay under three distinct transmission scenarios: varying delay taps ($L_{0}=6$, $L_{1}=3$, $L_{3}=4$), spatial distribution parameters ($R=15\,\text{m}$, $D_{3}=3\,\text{m}$), and Rician fading factors ($k_{\text{BR}}=3\,\text{dB}$, $k_{\text{RU}}=4\,\text{dB}$). The per-user arrival rate is fixed at $\lambda=9.5$, with all other parameters following the baseline in Section IV-A. Across all scenarios, the proposed algorithm consistently achieves the lowest delay, significantly outperforming benchmark methods. In contrast, the “max sum rate” approach leads to notably higher delays, while the “max minimum rate” shows moderate performance. The “random” strategy results in the worst delay across all settings. These results confirm the strong adaptability and delay minimization capability of the proposed algorithm under diverse transmission conditions. Its robustness against variations in different delay taps, user distribution, and fading characteristics highlights its potential for real-world deployment in dynamic wireless environments.
@@ V Conclusion
This paper introduced a novel DRL-based approach to minimize the average delay while ensuring user fairness in different traffic flows. Unlike traditional methods, our proposed solution extends the state information to encompass a comprehensive set of factors, including channel state information, the number of backlogged packets in buffer and the number of current arrival packets. To address the challenges posed by the hybrid action space, we develop a hybrid DRL framework, where the action space consists of both continuous and discrete components. Specifically, PPO-$\Theta$ is employed to optimize the continuous RIS reflection phase shifts, while PPO-$N$ is designed to handle the discrete subcarrier allocation decisions. Furthermore, to tackle the curse of dimensionality arising from multi-dimensional discrete variables, we introduced a multi-agent strategy to optimize subcarrier allocation decisions efficiently. Through extensive numerical evaluations, our method demonstrated significant improvements in reducing the average delay while maintaining user fairness, surpassing existing baseline methods. Additionally, the results highlighted the superior stability and adaptability of our approach. This work provides a robust and scalable solution to the delay optimization in RIS-assised OFDM systems, leveraging advanced machine learning techniques to optimize resource allocation in next-generation wireless networks, paving the way for more efficient and equitable network management.

### v4

@@ III-F Computational Complexity
The computational complexity of PPO-$\Theta$ is $\mathcal{O}(KNN_{t}(2M+1))$, while that of PPO-$N$ is $\mathcal{O}(KNN_{t})$. On the other hand, the active beamforming operation incurs a complexity of $\mathcal{O}(NMN_{t}+N)$. As a result, the overall computational complexity of the proposed algorithm is $\mathcal{O}(KNN_{t}(2M+1)+KNN_{t}+NMN_{t}+N)$. Since the max minimum rate benchmark shares the same overall algorithmic structure as the proposed method, its computational complexity is also $\mathcal{O}(KNN_{t}(2M+1)+KNN_{t}+NMN_{t}+N)$. Moreover, according to [37] and [38], the complexity of the max sum rate benchmark is $\mathcal{O}((MN)^{4.5}(KN_{t})^{1.5}+(K+1)^{4}+(K+1)^{2}KN_{t}N+NMN_{t}+N)$.
TABLE II: Hyper-Parameters
@@
PPO-$\Theta$ replay buffer size 1,000
PPO-N replay buffer size 1,000
@@ IV Numerical Results
In this section, a comprehensive evaluation of the performance of the proposed DRL-based algorithm is conducted. Various simulation scenarios and baseline comparisons are considered to demonstrate its effectiveness in optimizing the system performance.
@@ IV-A Simulation Settings
The packet capacity $L$ is set to 512 bits. In each episode, the communication duration is set to $T_{\text{sum}}=$ 1,000 time slots, with each time slot lasting 1 ms. Since the position of user remains nearly fixed within 1 second, we assume that the user positions are fixed throughout each episode. [^1] Note that, to accelerate the convergence of the algorithm, the user positions are randomly generated at each step within an episode during the training process. The packet arrival rate for each user is set to $\lambda=9.5$. Throughout this work, unless stated otherwise, identical arrival rates are assumed for all users, i.e., $\lambda=\lambda_{1}=\lambda_{2}=\cdots=\lambda_{K}$. The detailed hyperparameters are presented in Table II.
Figure 5: Return for each training episode with an increasing number of iterations.
@@ IV-B Baseline Methods
To reflect the performance of our proposed method, the following baseline methods are considered as follows:
  • Max sum rate: The algorithm proposed in [37] and [38] is respectively employed for subcarrier allocation and RIS reflection shift optimization, while MRT is used for beamforming at the BS.
  • Max minimum rate: This algorithm is the hybrid DRL algorithm with the reward to be $\min R_{k}[t]$, as described in Section III-E, and it has not undergone transfer learning.
  • Random: The RIS reflection shift and subcarrier allocation are set randomly. The active beamforming at the BS is designed using the MRT.
@@ IV-C Convergence and Validity Analysis
To validate the effectiveness and convergence of the proposed method, experiments are conducted. Figure 5 shows the cumulative reward for each training episode. The figure demonstrates a progressive increase in the cumulative reward throughout the training process, indicating the effectiveness of our proposed method. Besides, the performance gradually stabilizes in around 150 training episodes, despite some fluctuations due to channel variations and changes in the number of packets.
@@ IV-D Latency Performance Analysis
Figure 6 illustrates the average delay performance under varying packet arrival rates $\lambda$. The proposed algorithm consistently achieves the lowest average delay, demonstrating its superior capability in minimizing latency. In contrast, the “without RIS” scheme results in the highest delay, emphasizing the critical role of RIS in improving the delay performance. Although the “max sum rate” method performs better than baseline schemes, such as “random” and “without RIS”, its overall delay performance is clearly inferior to that of the proposed method. This is because it prioritizes the maximization of system throughput without considering user fairness or queue stability, which leads to higher delays for certain users. The “random” method yields persistently high delays, further underlining the importance of intelligent and delay-aware resource allocation. Moreover, as $\lambda$ increases, the number of packets arriving at the buffer within each time slot grows accordingly, leading to aggravated buffer backlog and, consequently, resulting in an incremental increase in the average delay.
(a) The backlogged packets in each buffer of the proposed method
(b) The backlogged packets in each buffer of the max minimum rate method
(c) Transmission rates of the proposed method
(d) Transmission rates of the max minimum rate method
Figure 8: The change of the backlogged packets and transmission rates in each buffer over time steps of the proposed method and max minimum rate method in brustiness scenarios.
(a) The backlogged packets in each buffer of the proposed method
(b) The backlogged packets in each buffer of the max minimum rate method
(c) Transmission rates of the proposed method
(d) Transmission rates of the max minimum rate method
Figure 9: The change of the backlogged packets and transmission rates in each buffer over time steps of the proposed method and max minimum rate method in different traffic flows scenarios.
Figure 10: Impact of the number of reflection elements $M$ on the delay performance.
@@ IV-E Latency Analysis in Brustiness Scenarios
To evaluate the proposed algorithm under bursty traffic, we simulate sudden data arrivals by injecting 300 packets into buffers 1, 2, and 3 at time slots 100, 400, and 700, respectively. Figure 7 presents a comparison of the delay performance between the proposed algorithm and the “max minimum rate” algorithm under different scenarios. It can be observed that, regardless of whether burstiness is present, the proposed algorithm consistently demonstrates superior performance, effectively validating its effectiveness in handling burstiness.
To further explain this advantage, we analyze a specific case by examining buffer backlogs and transmission rates during the first 300 time slots. As shown in Fig. 8LABEL:sub@Fig:packagePropsed, the proposed algorithm effectively reduces backlog by dynamically adjusting resource allocation in response to bursty arrivals. In contrast, Fig. 8LABEL:sub@Fig:packageMaxMin shows persistent congestion in buffer 3 under the “max minimum rate” approach, indicating inefficient allocation. Figure 8LABEL:sub@rateProposed further shows that the proposed method prioritizes users with heavier backlogs, rapidly increasing their transmission rates, such as user 3, leading to quick backlog clearance. Conversely, Fig. 8LABEL:sub@rateMaxMin reveals that the baseline allocates resources more rigidly, ignoring buffer states and failing to relieve congestion effectively, which results in higher delay.
@@ IV-F Impact of Reflection Elements $M$
Figure 10 illustrates the impact of the number of RIS reflection elements on the average delay performance under varying arrival rates $\lambda$. As observed, when $M=36$, the average delay increases sharply once $\lambda$ exceeds 9.0, suggesting that the system struggles to accommodate high traffic loads. With $M=64$, the delay grows more moderately but still experiences a considerable rise at elevated arrival rates. In contrast, when $M=100$, the delay remains consistently low across all values of $\lambda$, demonstrating that $M=100$ is sufficient enough to support the considered service load with $\lambda$ less than 10. This improvement is primarily attributed to the enhanced passive beamforming gain resulting from the increased number of RIS reflection elements, thereby improving the effective data rate. As a result, the system is able to accommodate higher traffic loads while maintaining low-latency communication performance.
@@ IV-G Latency Analysis under Different Traffic Flows
To further explain this advantage, we analyze a specific case by examining buffer backlogs and transmission rates during the first 300 time slots. In Fig. 9LABEL:sub@Fig:packagePropsed, the proposed algorithm keeps buffer backlogs balanced despite differing arrival rates, showing its fairness. Upon the burst at time slot 100, the algorithm quickly reallocates resources to clear buffer 3. In contrast, Fig. 9LABEL:sub@Fig:packageMaxMin shows that the “max minimum rate” approach leads to severe congestion in buffer 1 due to static rate-based allocation. Further, Fig. 9LABEL:sub@rateProposed reveals that the proposed method allocates more resources to high arrival rate users and dynamically adapts to queue states. Before time slot 100, user 3 has a lower rate due to low arrival, but receives rapid support when the burst occurs. However, Fig. 9LABEL:sub@rateMaxMin shows that the “max minimum rate” method fails to allocate resources and optimize beamforming based on buffer states, leading to increased delays.
TABLE III: The robustness of proposed approach in terms of average delay
@@
colspec = —X[c]—X[c]—X[c]—X[c]—X[c]—, hlines, vlines Algorithm & Delay taps ($L_{0}=6$, $L_{1}=3$, $L_{3}=4$) Distribution ($R=15$ m, $D_{3}=3$ m) Ricean factor ($k_{\mathrm{BR}}=3$ dB, $k_{\mathrm{RU}}=4$ dB)
@@ IV-H Jitter Analysis
$\frac{1}{K}\sum_{k=1}^{K}\sqrt{\frac{1}{m(k)}\sum_{g=1}^{m(k)}\bigg{(}T_{D}^{(k)}(g)-\frac{1}{m(k)}\sum_{g=1}^{m(k)}T_{D}^{(k)}(g)\bigg{)}^{2}}.$ (27)
As illustrated in Fig. 12, the proposed algorithm consistently achieves the lowest jitter across all values of $\lambda$, demonstrating its superior capability in balancing resource allocation and maintaining latency stability. In contrast, the “max-min rate” scheme yields slightly higher jitter, while the “max sum rate” approach, although effective in optimizing the overall throughput, results in significantly increased jitter. The “random” allocation and “without RIS” schemes exhibit the worst jitter performance, underscoring the effectiveness of both learning-based optimization and RIS assistance in enhancing system delay stability. Moreover, as the arrival rate $\lambda$ increases, jitter also exhibits a rising trend, indicating that the system becomes increasingly unstable under higher traffic loads and struggles to maintain consistent delay performance.
Figure 12: Jitter performance under varying packet arrival rates $\lambda$.
@@ IV-I Robustness Validation
Table III presents a robustness evaluation of the proposed algorithm in terms of average delay under three distinct transmission scenarios: varying delay taps ($L_{0}=6$, $L_{1}=3$, $L_{3}=4$), spatial distribution parameters ($R=15\,\text{m}$, $D_{3}=3\,\text{m}$), and Ricean fading factors ($k_{\text{BR}}=3\,\text{dB}$, $k_{\text{RU}}=4\,\text{dB}$). The per-user arrival rate is fixed at $\lambda=9.5$, with all other parameters following the baseline in Section IV-A. Across all scenarios, the proposed algorithm consistently achieves the lowest delay, significantly outperforming benchmark methods. In contrast, the “max sum rate” approach leads to notably higher delays, while the “max minimum rate” shows moderate performance. The “random” strategy results in the worst delay across all settings. These results confirm the strong adaptability and delay minimization capability of the proposed algorithm under diverse transmission conditions. Its robustness against variations in different delay taps, user distributions, and fading characteristics highlights its potential for real-world deployment in dynamic wireless environments.
@@ V Conclusion
This paper introduced a novel DRL-based approach to minimize the average delay while ensuring user fairness in different traffic flows. Unlike traditional methods, our proposed solution extends the state information to encompass a comprehensive set of factors, including CSI, the number of backlogged packets in buffer and the number of current arrival packets. To address the challenges posed by the hybrid action space, we developed a hybrid DRL framework, where the action space consists of both continuous and discrete components. Specifically, PPO-$\Theta$ was employed to optimize the continuous RIS reflection phase shifts, while PPO-$N$ was designed to handle the discrete subcarrier allocation decisions. Furthermore, to tackle the curse of dimensionality arising from multi-dimensional discrete variables, we introduced a multi-agent strategy to optimize subcarrier allocation decisions efficiently. Through extensive numerical evaluations, our method demonstrated significant improvements in reducing the average delay while maintaining user fairness, surpassing existing baseline methods. Additionally, the results highlighted the superior stability and adaptability of our approach. This work provides a robust and scalable solution to the delay minimization in RIS-assised OFDM systems, leveraging advanced machine learning techniques to optimize resource allocation in next-generation wireless networks, paving the way for more efficient and equitable network management.